COMPUTATION OF COLLISION-INDUCED ABSORPTION BY SIMPLE MOLECULAR COMPLEXES, FOR ASTROPHYSICAL APPLICATIONS

\underline{MARTIN ABEL}, LOTHAR FROMMHOLD, Department of Physics, The University of Texas at Austin, Austin, TX 78712; XIAOPING LI, KATHARINE L. C. HUNT, Department of Chemistry, Michigan State University, East Lansing, MI 48824.

The interaction-induced absorption by collisional pairs of H_{2} molecules is an important opacity source in the atmospheres of various types of planets and cool stars, such as late stars, low-mass stars, brown dwarfs, cool white dwarf stars, the ambers of the smaller, burnt out main sequence stars, exoplanets, etc., and therefore of special astronomical interest ^{a}. The emission spectra of cool white dwarf stars differ significantly in the infrared from the expected blackbody spectra of their cores, which is largely due to absorption by collisional H_{2}–H_{2}, H_{2}–He, and H_{2}–H complexes in the stellar atmospheres. Using quantum-chemical methods we compute the atmospheric absorption from hundreds to thousands of kelvin b. Laboratory measurements of interaction-induced absorption spectra by H_{2} pairs exist only at room temperature and below. We show that our results reproduce these measurements closely ^{c}, so that our computational data permit reliable modeling of stellar atmosphere opacities even for the higher temperatures ^{d}. First results for H_{2}–He complexes ^{e} have already been applied to astrophysical models ^{f} and have shown great improvements in these models.

aL. Frommhold, Collision-Induced Absorption in Gases, Cambridge University Press, Cambridge, New York, 1993 and 2006

bX. Li, K. L. C. Hunt, F. Wang, M. Abel, and L. Frommhold, Collision-Induced Infrared Absorption by Molecular Hydrogen Pairs at Thousands of Kelvin, Int. J. of Spect., vol. 2010, Article ID 371201, 11 pages, 2010. doi: 10.1155/2010/371201

cM. Abel, L. Frommhold, X. Li, and K. L. C. Hunt, Collision-induced absorption by H_{2} pairs: From hundreds to thousands of Kelvin, J. Phys. Chem. A, 115, 6805-6812, 2011

dL. Frommhold, M. Abel, F. Wang, M. Gustafsson, X. Li, and K. L. C. Hunt, ”Infrared atmospheric emission and absorption by simple molecular complexes, from first principles”, Mol. Phys. 108, 2265, 2010

eM. Abel, L. Frommhold, X. Li, and K. L. C. Hunt, Infrared absorption by collisional H_{2}–He complexes at temperatures up to 9000 K and frequencies from 0 to 20000 cm^{−1}, J. Chem. Phys., 136, 044319, 2012

fD. Saumon, M. S. Marley, M. Abel, L. Frommhold, and R. S. Freedman, New H_{2} collision-induced absorption and NH_{3} opacity and the spectra of the coolest brown dwarfs, Astrophysical Journal, 2012